Carbon dioxide is librated in the reaction of the ore hematite (fe²o³)with coke (c) for the production of iron.if the hypothetically as a member of a blast furnace you are told you have 20.5Mg (megagrams)of Fe2O³and 2.84Mg of coke hand.
a.which would you order first- another shipment of iron ore or coke? justify
b.how many megagrams of iron you make with the materials u have (Mg =megagrams)
1
Expert's answer
2020-02-17T07:09:56-0500

2Fe2O3 + 3C = 4Fe + 3CO2

We have:

n(Fe2O3) = 20.5/160 = 0.128 Mmol

n(C) = 2.84/12 = 0.237 Mmol

n(C)/n(Fe2O3) = 1.84

Coke is in excess, that is why I would order hematite first.

m(Fe) = 0.128 * 2 * 55,8 = 14.28 Mg